# OVERVIEW ARTICLE

### v1

@@ 3.2   Legal Issues Arising from Memorization
The above described phenomenon of memorization introduces two primary legal risks discussed in the following. [^23] For a concise analysis of the following issues in English see also Dornis, 2025a .
@@ 3.2.1   Sharing models that contain (partially) memorized training data may infringe copyright
From a legal perspective, it does not matter whether we understand the internal code or not. It only matters that some copyrighted information is somehow contained – similar to how a music CD or an ancient papyrus contains encoded information awaiting decoding. Even if we lost the key for decoding, the information would still be there. Consequently, sharing a model that has partly memorized some copyrighted material – even if this was lawfully obtained – may infringe copyright.
@@ 3.2.2   Generating output containing (partially) memorized training data may infringe copyright
Even when models are not shared, their outputs may still infringe copyright if they contain memorized material. Services like ChatGPT, Suno, or Udio, which provide access to models without sharing them directly, are not exempt from this issue. If a model reproduces copyrighted content upon request, the service operator may be held liable under the copyright laws of the jurisdiction where the output was generated – i.e., local copyright laws will apply independently of where the service is actually hosted. Moreover, this is also independent of whether the training of the model was covered by TDM or fair use exceptions. This results in global compliance challenges for operators who must navigate differing legal standards across countries. At the same time, this allows rightsholders to litigate against powerful global AI tech corporations in local courts. For instance, the German society for musical performing and mechanical reproduction rights (GEMA) has already started lawsuits against OpenAI in November 2024 and Suno AI in January 2025 claiming the unlicensed reproduction of song texts and music respectively that is uncannily similar to copyrighted material of German artists.
@@ 3.3   Conclusion
Memorization in generative AI models exemplifies the complex interplay between technological capabilities and legal frameworks. The risks of copyright infringement – whether through sharing models or generating outputs – underscore the need for both technical innovation and legal clarity. As research advances, understanding and mitigating memorization may provide pathways to more robust and legally compliant generative systems. Meanwhile, fostering dialogue between researchers, policymakers, and rights holders remains essential to navigating this intricate legal terrain.
@@ 4.   Towards Fair and Pragmatic Documentation of Training Data Sources
A critical yet often overlooked aspect of generative AI model development is the documentation of training data sources. While much of the debate surrounding AI training data focuses on legality and ethical concerns, there is also a pragmatic dimension: the transparent and structured documentation of data provenance. Recent legislative developments, such as the European Union’s AI Act, emphasize the need for such documentation. Article 53 of the AI Act requires providers of general-purpose AI models to “draw up and make publicly available a sufficiently detailed summary about the content used for training”. [^25] European Union, “Regulation (EU) 2024/1689 of the European Parliament and of the Council on the AI Act,” 2024. Available: https://eur-lex.europa.eu/eli/reg/2024/1689/oj However, the notion of a “sufficiently detailed summary” remains vague and open to interpretation, leading to concerns about its practical implementation.
@@ 4.1   Challenges in Documenting Training Data
From a technical standpoint, large-scale data collection and curation often involve automated scraping, database aggregation, and third-party dataset integration. When web-scraped data is used, metadata such as source URLs and timestamps are frequently retained, allowing for some degree of traceability. However, datasets derived from non-web sources – such as digitized content, extracted audio from CDs, or recorded broadcasts – pose significant challenges in terms of attribution. In many cases, such datasets lack inherent metadata linking them to original rightsholders, requiring additional effort to establish provenance.
Given these challenges, there is a risk that compliance with transparency requirements will be implemented in a superficial manner, providing only vague descriptions of dataset composition without meaningful traceability. This approach does little to satisfy the legitimate interests of rightsholders, researchers, or policymakers seeking to understand the origins and licensing conditions of the training data.
@@ 4.2   Leveraging Music Information Retrieval for Dataset Attribution
The field of MIR offers a set of tools that can contribute significantly to improving dataset documentation. Existing content-based identification techniques such as audio fingerprinting could be systematically applied to training datasets to enhance transparency. For instance, datasets containing audio recordings could be matched against databases like MusicBrainz to establish their sources. Similarly, MIDI datasets, such as the Lakh MIDI dataset (Raffel,, 2016), have already been partially linked to known recordings via cross-referencing with the Million Song Dataset (Bertin-Mahieux et al.,, 2011).
@@ 4.3   Pragmatic Approaches to Documentation
To balance the needs of rightsholders and the practical constraints of AI developers, a tiered approach to documentation can be adopted:
@@ 4.4   Benefits Beyond Legal Compliance
The ISMIR community is uniquely positioned to take the lead in establishing best practices for dataset documentation in generative AI research. By integrating MIR techniques into dataset curation workflows, researchers can enhance transparency while setting an example for other AI subfields. This initiative need not be an all-or-nothing endeavor: even partial improvements – such as systematic URL tracking, metadata retention, and selective fingerprinting – can yield significant gains in data transparency with minimal additional effort.
@@ 4.5   Recommendations for Action
Promote Collaboration Between Stakeholders: Encourage collaboration between AI developers, rightsholders, and researchers to establish shared standards for dataset documentation. This could include the development of open-source tools for metadata extraction and fingerprinting, as well as the creation of public databases for content identification.
@@
By embedding principles of fair attribution into dataset creation and usage, the ISMIR community can set a precedent for responsible AI development. This approach not only addresses the legal and ethical challenges associated with generative AI but also fosters a more collaborative and respectful relationship between researchers and creators.
@@ 5.   Conclusions and Outlook
The rapid advancement of generative AI, particularly in the domain of music generation, has brought to the forefront a complex interplay between technological innovation, legal frameworks, and ethical considerations. This paper has explored the multifaceted challenges associated with training generative AI models, focusing on the legal implications of data collection, the phenomenon of training data memorization, and the need for transparent and fair documentation of training data sources. Below, we summarize the key findings and provide an outlook on future directions for research and policy.
5.1   Summary of Key Findings
Legal Frameworks for Training Data Collection: The legal landscape for generative AI training data remains fragmented and uncertain. While the EU’s Text and Data Mining (TDM) exceptions and the US fair use doctrine provide some guidance, it is far from clear whether they can provide a valid defense against claims of copyright infringement that may occur during GenAI training. Our analysis reveals that generative AI training fundamentally diverges from traditional TDM activities, as it involves the creation of derivative works rather than the extraction of knowledge. This raises significant questions about the applicability of existing legal exceptions, particularly in jurisdictions where no explicit provisions for AI training exist.
Memorization and Its Legal Implications: Generative AI models have been shown to memorize substantial portions of their training data, even without explicit storage mechanisms. This phenomenon introduces two primary legal risks: (1) sharing models that contain memorized data may infringe copyright, and (2) generating outputs that reproduce memorized content may also lead to infringement. These risks are exacerbated by the global nature of AI services, which must navigate differing legal standards across jurisdictions. Technical mitigation strategies, such as input filtering and improved model design, remain in their infancy and cannot yet fully eliminate these risks.
Fair and Pragmatic Documentation of Training Data: Transparent documentation of training data sources is essential for ensuring legal compliance, fostering trust, and respecting the rights of creators. While legislative efforts like the EU AI Act mandate the disclosure of training data summaries, the practical implementation of such requirements remains challenging. Leveraging MIR techniques, such as audio fingerprinting and metadata matching, can significantly enhance dataset transparency. We proposed tiered approach to documentation – ranging from basic metadata to advanced fingerprinting – that offers a pragmatic solution that balances the needs of rightsholders with the constraints of AI developers.
@@ 5.2   Implications for the MIR Community
The findings of this paper have several implications for the Music Information Retrieval (MIR) community:
Legal Awareness: Researchers and practitioners must remain vigilant about the legal risks associated with generative AI training – particularly when reusing datasets originally created for TDM tasks. Explicit permission from rightsholders may be required for new uses of existing datasets.
@@
Ethical Responsibility: By adopting best practices for dataset documentation and attribution, the MIR community can demonstrate its commitment to ethical AI development and foster a more collaborative relationship with creators and rightsholders.
@@ 5.3   Outlook and Future Directions
Harmonization of Legal Frameworks: The global nature of AI research and data use underscores the need for international dialogue on harmonizing legal frameworks for generative AI training. Policymakers, researchers, and industry stakeholders must collaborate to develop clear and consistent guidelines that balance innovation with the protection of intellectual property rights. This paper should have made it clear that such guidelines can have a great impact on MIR research. We therefore hope that this paper will ecourage the MIR community to get more involved in the ongoing debate.
@@
Engagement with Creators and Rightsholders: Building trust and collaboration with creators and rightsholders is critical for the sustainable development of generative AI. Initiatives such as public fingerprint databases and machine-readable opt-out mechanisms can empower creators to assert their rights while enabling researchers to comply with legal and ethical standards. This could be supported by technical solutions developed within the MIR community.
@@ 5.4   Final Thoughts
The intersection of generative AI and copyright law presents both challenges and opportunities for the MIR community. By addressing the legal, technical, and ethical dimensions of AI training, researchers can contribute to the development of generative systems that are not only innovative but also respectful of the rights and contributions of creators. Ultimately, this approach will help to sustain human creators and thereby guarantee the future of human creativity as the basis for generative AI – a goal that everybody should be able to subscribe to. The path forward requires a collaborative effort, involving researchers, policymakers, rightsholders, and industry stakeholders, to ensure that generative AI serves as a tool for creativity and progress rather than a source of conflict and uncertainty.
In conclusion, while the road ahead is complex, the MIR community is well-equipped to navigate these challenges and lead the way toward a future where generative AI and copyright law coexist harmoniously. By embracing transparency, innovation, and ethical responsibility, we can unlock the full potential of generative AI while upholding the principles of fairness and respect that underpin creative expression. As ISMIR community, we can demonstrate that we deeply care about the art of music and the artists by making an effort to properly acknowledge the people whose data we use and giving them a chance to provide or withdraw consent.
@@ 6.   Competing interests
Parts of this article are based on an interdisciplinary study on generative AI and copyright law in Germany and the EU conducted by the authors in 2024 and funded by the German Authors‘ Rights Initiative (Initiative Urheberrecht). The contents of the resulting report (Dornis and Stober,, 2024) as well as this article exclusively reflect the views and assessments of the authors.

### v2

@@ 3.3  Legal Issues Arising from Memorization
The above described phenomenon of memorization introduces two primary legal risks discussed in the following. [^30] For a concise analysis of the following issues in English see also Dornis (2025a).
@@ 3.3.1  Sharing models that contain (partially) memorized training data may infringe copyright
From a legal perspective, it does not matter whether we understand the internal code or not. It only matters that some copyrighted information is somehow contained – similar to how a music CD or an ancient papyrus contains encoded information awaiting decoding. Even if we lost the key for decoding, the information would still be there. Consequently, reproducing and copying, e.g. by sharing a model that has partly memorized some copyrighted material – even if this was lawfully obtained – may infringe copyright.
The geometric perspective of Ross and other authors (2024) also illuminates why detecting memorization is challenging: a model may contain internal representations of copyrighted works without ever reproducing them, if those works lie on the learned manifold but are never triggered by an appropriate prompt. This aligns with our earlier logical distinction between memorization (internal storage) and regurgitation (observable output). The former is necessary but not sufficient for the latter, and the absence of observable regurgitation does not prove the absence of memorization.
Whether sharing a model constitutes copyright infringement depends on what the model contains and how it is used. If a model is shared for research purposes and contains memorized snippets, the legal outcome may differ from a scenario where the model is commercialized. Courts may consider the purpose of sharing, the likelihood of extraction, and the potential market harm. However, the mere fact that a model embodies copyrighted material without authorization creates a risk, especially if the model is made publicly available. [^32] For this aspect see Regional Court Munich I, final judgment of 11 November 2025, docket# 42 O 14139/24, GRUR-RS 2025, 30204 – ChatGPT.https://www.gesetze-bayern.de/Content/Document/Y-300-Z-GRURRS-B-2025-N-30204.
@@ 3.3.2  Generating output containing (partially) memorized training data may infringe copyright
Even when models are not shared, their outputs may still infringe copyright if they contain memorized material. Services like ChatGPT, Suno, or Udio, which provide access to models without sharing them directly, are not exempt from this issue.32 If a model reproduces copyrighted content upon request, the service operator may be held liable under the copyright laws of the jurisdiction where the output was generated – i.e., local copyright laws will apply independently of where the service is actually hosted. Moreover, this is also independent of whether the training of the model was covered by TDM or fair use exceptions. This results in global compliance challenges for operators who must navigate differing legal standards across countries. At the same time, this allows rightholders to litigate against powerful global AI tech corporations in local courts.
As already mentioned, courts have already decided. In particular in Germany, the Munich I Regional Court found copyright infringement in GenAI training. The German society for musical performing and mechanical reproduction rights (GEMA) filed lawsuits against OpenAI in November 2024 and Suno AI in January 2025 claiming the unlicensed reproduction of song texts and music respectively that is uncannily similar to copyrighted material of German artists. In November 2025, the Munich I Regional Court ruled in GEMA’s favor, holding that the memorization of song lyrics in OpenAI’s GPT-4 and GPT-4o models constituted unlawful reproduction under German copyright law, rejecting OpenAI’s text and data mining defense, and placing responsibility on OpenAI rather than its users. [^33] Regional Court Munich I, final judgment of 11 November 2025, docket# 42 O 14139/24, GRUR-RS 2025, 30204 – ChatGPT. OpenAI has appealed the decision, which is now pending before the Munich Higher Regional Court. The lawsuit against Suno AI is still ongoing. More lawsuits must be expected. In addition to the German case, there is litigation in Italy and Hungary – the later case now pending before the Court of Justice of the European Union (CJEU). [^34] CJEU, Case C-250/25, 3 April 2025. (Referring court: Budapest Környéki Törvényszék, Hungary) https://infocuria.curia.europa.eu/tabs/document/C/2025/C-0250-25-00000000RP-01-P-01/DDP/300681-EN-1-pdf.
Notably, although generating an output that reproduces a copyrighted work may principally infringe, here as well context matters. If a researcher generates a sample for non-commercial analysis, it might, for instance, be considered fair use in the U.S., whereas a company selling that output would face liability. In general and in practically all copyright regimes, the threshold for infringement is whether the output is substantially similar to the original and whether it harms the market.
3.3.3  Potential Mitigation Strategies
@@
Furthermore, Ghosh and other authors (2025) caution that improved learning can have opposing effects on different types of memorization: it may decrease contextual memorization while simultaneously increasing recollection-based memorization. This suggests that mitigation strategies must be carefully evaluated against multiple metrics to ensure they do not inadvertently exacerbate other forms of memorization risk.
@@ 3.4  Conclusion
Memorization in generative AI models exemplifies the complex interplay between technological capabilities and legal frameworks. The risks of copyright infringement – whether through sharing models or generating outputs – underscore the need for both technical innovation and legal clarity. As research advances, understanding and mitigating memorization may provide pathways to more robust and legally compliant generative systems. Meanwhile, fostering dialogue between researchers, policymakers, and rightholders remains essential to navigating this intricate legal terrain.
@@ 4.  Towards Fair and Pragmatic Documentation of Training Data Sources
The preceding sections have established two fundamental challenges for generative AI training under current copyright law. First, as argued in Section 2, the training of generative models may fall outside the scope of existing TDM exceptions and fair use doctrines. This means that any use of copyrighted material for training may require explicit permission from rightholders – a requirement that appears impossible to satisfy for models trained on vast, web-scale datasets. Second, as shown in Section 3, even if training itself could somehow be justified, the phenomenon of memorization creates independent legal risks: models may reproduce copyrighted content in their outputs, and sharing the models themselves may constitute distribution of infringing copies.
These two challenges converge on a single point: the need for transparency about what data was used to train a model. Without such transparency, rightholders cannot know whether their works have been used, cannot assert their opt-out rights where they exist, and cannot seek redress for infringing reproductions. Without transparency, AI developers cannot credibly demonstrate that they have made good-faith efforts to respect copyright, cannot identify and mitigate memorization risks, and cannot defend themselves against infringement claims. And without transparency, the research community cannot study the relationship between training data composition and model behavior—including memorization—in a systematic way.
This section therefore addresses a question that is both practical and legally essential: How can training data sources be documented in a way that satisfies the needs of rightholders, researchers, and AI developers alike? We do not propose that documentation alone solves all legal problems, but we argue that it is an indispensable first step toward any workable solution.
This need for transparency is already reflected in recent legislation. Article 53 of the European Union’s AI Act, for instance, requires providers of general-purpose AI models to “draw up and make publicly available a sufficiently detailed summary about the content used for training”. [^35] European Union, “Regulation (EU) 2024/1689 of the European Parliament and of the Council on the AI Act,” 2024. https://eur-lex.europa.eu/eli/reg/2024/1689/oj. However, the notion of a “sufficiently detailed summary” remains vague and open to interpretation, leading to concerns about its practical implementation.
These concerns are not merely theoretical. The Open Source AI Definition (OSAID 1.0) recently adopted a nearly identical requirement for open models, mandating a “sufficiently detailed summary” of training data. [^36] Open Source Initiative, “Open Source AI Definition 1.0,” 2024. https://opensource.org/ai/open-source-ai-definition. Critics have already labeled this as “open-washing,” arguing that without access to the actual data, such summaries provide little meaningful accountability. [^37] Open Future and Mozilla Foundation, “Sufficiently detailed summary: v2.0 of the blueprint for GPAI training data,” 2024. https://openfuture.eu/blog/sufficiently-detailed-summary-v2-0-of-the-blueprint-for-gpai-training-data/. This serves as a cautionary tale: if the AI Act’s requirement is implemented without clear standards, it may similarly fail to deliver the transparency that rightholders and researchers need. Practical guidance, such as the blueprint developed by Open Future and the Mozilla Foundation, offers a path forward by specifying what a sufficiently detailed summary could concretely entail.37
The remainder of this section examines the technical challenges of documentation, the tools that MIR research can contribute, and a pragmatic tiered approach that balances feasibility with accountability.
@@ 4.1  Challenges in Documenting Training Data
From a technical standpoint, large-scale data collection and curation often involve automated scraping, database aggregation, and third-party dataset integration. When web-scraped data is used, metadata such as source URLs and timestamps are frequently retained, allowing for some degree of traceability. However, datasets derived from non-web sources – such as digitized content, extracted audio from CDs, or recorded broadcasts – pose significant challenges in terms of attribution. In many cases, such datasets lack inherent metadata linking them to original rightholders, requiring additional effort to establish provenance.
Given these challenges, there is a risk that compliance with transparency requirements will be implemented in a superficial manner, providing only vague descriptions of dataset composition without meaningful traceability. This approach does little to satisfy the legitimate interests of rightholders, researchers, or policymakers seeking to understand the origins and licensing conditions of the training data.
@@ 4.2  Leveraging Music Information Retrieval for Dataset Attribution
The field of MIR offers a set of tools that can contribute significantly to improving dataset documentation. Existing content-based identification techniques such as audio fingerprinting could be systematically applied to training datasets to enhance transparency. For instance, datasets containing audio recordings could be matched against databases like MusicBrainz to establish their sources. Similarly, MIDI datasets, such as the Lakh MIDI dataset (Raffel, 2016), have already been partially linked to known recordings via cross-referencing with the Million Song Dataset (Bertin-Mahieux et al., 2011).
@@ 4.3  Pragmatic Approaches to Documentation
To balance the needs of rightholders and the practical constraints of AI developers, a tiered approach to documentation can be adopted:
@@
To illustrate the tiered approach, consider popular MIR datasets. GTZAN, a collection of 1000 audio clips, often lacks detailed provenance. Sturm’s comprehensive analysis (including audio fingerprinting) documents repetitions, mislabelings, and distortions that challenge its use without careful curation (Sturm, 2013).
MusDB18, a dataset of 150 full-track songs with isolated stems, provides basic track metadata such as artist and song names (Rafii et al., 2017). While this enables basic attribution, it lacks the unique identifiers (e.g., ISRCs) that would allow linking to rightholders’ databases.
The MTG-Jamendo dataset (Bogdanov et al., 2019) exemplifies the advanced documentation tier. It comprises over 55,000 tracks from the Jamendo platform under explicit Creative Commons licenses, accompanied by rich metadata. The licensing status is known upfront, and fingerprinting techniques can be applied to verify attribution and enable rightholder opt-out mechanisms.
User-generated content datasets like AudioSet (Gemmeke et al., 2017), a large-scale collection of over two million human-labeled 10-second sound clips drawn from YouTube videos, exemplify the most challenging scenario. These datasets rely on platform-specific identifiers (YouTube video IDs) that may not persist, [^38] As noted in the dataset documentation, when videos are removed from YouTube, the corresponding audio clips become unavailable; a 2023 re-download of the dataset could only retrieve approximately 84% of the original balanced training set and 84% of the evaluation set. See https://huggingface.co/datasets/agkphysics/AudioSet. and the content itself lacks inherent, rights-holder-verified metadata. For such datasets, thorough documentation should include the original video IDs and timestamps, and researchers could additionally employ fingerprinting techniques to help maintain traceability despite the instability of the source platform.
Recent work at ISMIR, such as by Evans et al. (2024) and Chen et al. (2024b), shows varying degrees of dataset transparency, underscoring the need for community-wide standards.
@@ 4.4  Benefits Beyond Legal Compliance
The MIR community is uniquely positioned to take the lead in establishing best practices for dataset documentation in generative AI research. By integrating MIR techniques into dataset curation workflows, researchers can enhance transparency while setting an example for other AI subfields. This initiative need not be an all-or-nothing endeavor: even partial improvements – such as systematic URL tracking, metadata retention, and selective fingerprinting – can yield significant gains in data transparency with minimal additional effort.
@@ 4.5  Recommendations for Action
Promote Collaboration Between Stakeholders: Encourage collaboration between AI developers, rightholders, and researchers to establish shared standards for dataset documentation. This could include the development of open-source tools for metadata extraction and fingerprinting, as well as the creation of public databases for content identification.
@@
By embedding principles of fair attribution into dataset creation and usage, the MIR community can set a precedent for responsible AI development. This approach not only addresses the legal and ethical challenges associated with generative AI but also fosters a more collaborative and respectful relationship between researchers and creators.
@@ 5.  Conclusions and Outlook
The rapid advancement of generative AI, particularly in the domain of music generation, has brought to the forefront a complex interplay between technological innovation, legal frameworks, and ethical considerations. This paper has explored the multifaceted challenges associated with training generative AI models, focusing on the legal implications of data collection, the phenomenon of training data memorization, and the need for transparent and fair documentation of training data sources. We have argued that existing TDM and fair use exceptions are ill-suited for GenAI training, that memorization introduces substantial legal risks, and that improved documentation practices can help mitigate those risks while fostering trust.
@@ 5.1  Implications for the MIR Community
The findings of this paper have several implications for the MIR community:
Legal Awareness: Researchers and practitioners must remain vigilant about the legal risks associated with generative AI training – particularly when reusing datasets originally created for TDM tasks. Explicit permission from rightholders may be required for new uses of existing datasets.
@@
Ethical Responsibility: By adopting best practices for dataset documentation and attribution, the MIR community can demonstrate its commitment to ethical AI development and foster a more collaborative relationship with creators and rightholders. [^39] For a philosophical perspective on the exploitative nature of AI art, see Goetze (2024).
@@ 5.2  Outlook and Future Directions
Harmonization of Legal Frameworks: The global nature of AI research and data use underscores the need for international dialogue on harmonizing legal frameworks for generative AI training. Policymakers, researchers, and industry stakeholders must collaborate to develop clear and consistent guidelines that balance innovation with the protection of intellectual property rights. This paper should have made it clear that such guidelines can have a great impact on MIR research. We therefore hope that this paper will encourage the MIR community to get more involved in the ongoing debate.
@@
Engagement with Creators and Rightholders: Building trust and collaboration with creators and rightholders is critical for the sustainable development of generative AI. Initiatives such as public fingerprint databases and machine-readable opt-out mechanisms can empower creators to assert their rights while enabling researchers to comply with legal and ethical standards. This could be supported by technical solutions developed within the MIR community.
Community Action: Given that the most egregious copyright violations often occur in commercial AI systems, the MIR community faces a choice: Should we accept sponsorship from companies that engage in massive infringement? Should we ban papers that use datasets lacking proper documentation? These are difficult questions, but they deserve open debate. By establishing clear ethical guidelines, the community can signal that it values integrity and will not tacitly endorse harmful practices.
@@ 5.3  Final Thoughts
The intersection of generative AI and copyright law presents both challenges and opportunities for the MIR community. By addressing the legal, technical, and ethical dimensions of AI training, researchers can contribute to the development of generative systems that are not only innovative but also respectful of the rights and contributions of creators. Ultimately, this approach will help to sustain human creators and thereby guarantee the future of human creativity as the basis for generative AI – a goal that everybody should be able to subscribe to. The path forward requires a collaborative effort, involving researchers, policymakers, rightholders, and industry stakeholders, to ensure that generative AI serves as a tool for creativity and progress rather than a source of conflict and uncertainty.
In conclusion, while the road ahead is complex, the MIR community is well-equipped to navigate these challenges and lead the way toward a future where generative AI and creativity coexist harmoniously. By embracing transparency, innovation, and ethical responsibility, we can unlock the full potential of generative AI while upholding the principles of fairness and respect that underpin creative expression. As the MIR community, we can demonstrate that we deeply care about the art of music and the artists by making an effort to properly acknowledge the people whose data we use and giving them a chance to provide or withdraw consent.
@@ 6.  Competing interests
Parts of this article are based on an interdisciplinary study on generative AI and copyright law in Germany and the EU conducted by the authors in 2024 and funded by the German Authors’ Rights Initiative (Initiative Urheberrecht). The contents of the resulting report (Dornis and Stober, 2024) as well as this article exclusively reflect the views and assessments of the authors.